OpenStudy (anonymous):

simplify: 9(a^-2)b over (b^-1)((8a)^-1/3))

OpenStudy (anonymous):

can't really make sense of this

OpenStudy (darthsid):

\[\frac{9a^{-2}b}{b^{-1}(8a)^{-\frac{1}{3}}}\] is this it?

OpenStudy (anonymous):

yeah its this: \[(9a ^{-2}b) \div (b ^{-1}\sqrt[3]{8a})\]

OpenStudy (anonymous):

when i tried to simplify it i got \[9b ^{2} \div (2a ^{7/3})\]

OpenStudy (anonymous):

\[\frac{9 (8a)^{\frac{1}{3}} a^{(-2)}b^2}{1}\]

OpenStudy (anonymous):

\[\frac{9 *2 (a)^{\frac{1}{3}} a^{(-2)}b^2}{1}\] \[\frac{18 (a)^{-\frac{5}{3}}b^2}{1}\]

OpenStudy (anonymous):

There are two different problems here. imranmeah has a negative 1/3 on the bottom which brings it to the top...but sunyah has a cube root on the bottom (which would stay there). Is it \[\frac{9a^{-2}b}{b^{-1}(8a)^{1/3}}\]or\[\frac{9a^{-2}b}{b^{-1}(8a)^{-1/3}}\]?

OpenStudy (anonymous):

it is just a cube root- that can be changed to (1/3) but not NEGATIVE... oohhh i see my mistake. :) THANK YOU!

OpenStudy (anonymous):

Then your answer of \[\frac{9b^{2}}{2a^{7/3}}\] would be right.

OpenStudy (anonymous):

thank youu i must have typed it wrong into wolfram alpha to check, got imranmeah91's answer, and was super confused because my work on paper seemed right. :)

OpenStudy (anonymous):

that's why i always trust my own pencil :)